The sum of two consecutive odd integers is 21 less than three times the larger integer. Find the numbers.
Answer by ankor@dixie-net.com(22740)

The sum of two consecutive odd integers is 21 less than three times the larger integer.
Two number n, (n-2)
n + (n-2) = 3n - 21
2n - 3n = -21 + 2
-n = -19
n = 19
:
Find the numbers. 17 and 19